a photon collide inelastically to a ground state H- atom. Colliding photon has 10.2 ev energy. in time difference of microsecond range an another photon with 15 ev energy collide with same H-atom. What will be recorded by the detector.

a. two photons of 10.2 ev energy
b.two photon of 1.4 ev energy
c. one photon with 10.2 ev energy and one electron with 1.4 ev energy
d. one photon with 10.2 ev energy another photon with 1.4 ev energy
Answer #1


nome 13 A Photon of Energy = 10.2er isev 10.2ev £, --13.6 = -13.6ev Ez = -13 6 = -13-6 E = -3.4er HT E2 - E = (-3.4 CV) - (-1